OpenStudy (anonymous):

Simplify (x – 1)(x2+ 5x + 2) x3+ 6x2+ 7x – 2 x3+ 4x2– 3x – 2 x3– x2+ 2x – 2 x3+ 5x2– 6x – 2

OpenStudy (anonymous):

\[(x – 1)(x^2+ 5x + 2)=(x^3+5x^2+2x)-(x^2+5x+2)= x^3+4x^2-3x-2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

We just multiply \[(x2+ 5x + 2)\] by x then by 1 and subtract. See: http://www.intmath.com/basic-algebra/1-addition-subtraction-algebra.php

OpenStudy (anonymous):